# TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning

### v1

@@ Abstract
Large Language Models (LLMs) present significant computational and memory challenges due to their extensive size, making pruning essential for their efficient deployment. Existing one-shot pruning methods often apply uniform sparsity constraints across layers or within each layer, resulting in suboptimal performance, especially at high sparsity ratios. This work introduces TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that applies varying sparsity ratios to individual output dimensions (rows) within each layer. TRIM employs an iterative adjustment process guided by quality metrics to optimize dimension-wise sparsity allocation, focusing on reducing variance in quality retention across outputs to preserve critical information. TRIM can be seamlessly integrated with existing layer-wise pruning strategies. Our evaluations on perplexity and zero-shot tasks across diverse LLM families (Qwen2.5, LLaMA-2, and OPT) and sparsity levels demonstrate that TRIM achieves new state-of-the-art results and enhances stability. For instance, at 80% sparsity, TRIM reduces perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to baseline methods. We conclude that fine-grained, dimension-wise sparsity adaptation is crucial for pushing the limits of extreme LLM compression. Code available at: https://github.com/flobk/TRIM
@@ 1 Introduction
Large language models (LLMs) have gained prominence because of their generalized problem-solving abilities (Brown et al., 2020; OpenAI et al., 2024; Zhao et al., 2025). A significant driver of this success has been the exponential scaling of model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Increasing the number of parameters has improved downstream performance and enabled emergent behaviors (Wei et al., 2022), such as in-context learning (Brown et al., 2020) and reasoning (Yao et al., 2023; Wei et al., 2023). However, this growth in model size leads to substantial memory and computational demands, which pose considerable deployment challenges (Zhou et al., 2024).
@@
As a result, one-shot pruning methods that reduce LLM parameters without retraining to compute weight masks have gained prominence (Han et al., 2015; Frantar and Alistarh, 2023; Sun et al., 2024). A challenge with current approaches is that they often impose predefined structural constraints, such as uniform sparsity across all network layers or outputs. This imposed structure is arbitrary and can be detrimental to pruning performance, particularly as LLMs exhibit unique weight and activation characteristics, such as prominent outlier features and highly skewed activation distributions (Kovaleva et al., 2021; Dettmers et al., 2022; Rudman et al., 2023).
Layer-wise sparsity adaptation methods aim to mitigate this by assigning distinct sparsity ratios to different layers of the model, determined through algorithmic search (Li et al., 2024; Xu et al., 2025), heuristic-based approaches (Yin et al., 2024; Sun et al., 2025), or more theoretically grounded strategies (Lu et al., 2024). While this offers more flexibility than globally uniform sparsity, these methods still apply a consistent sparsity target within each layer and can struggle under extreme sparsity levels, such as removing more than 70% of model parameters.
In this work, we propose TRIM: (Targeted Row-wise Iterative Metric-driven pruning). TRIM builds on the principles of one-shot pruning algorithms like Wanda (Sun et al., 2024), which compare and remove weights at each layer iteratively. Current methods apply a uniform sparsity constraint across all output dimensions, however, TRIM introduces a more granular approach. TRIM calculates unique sparsity ratios for individual output dimensions (corresponding to a row in the layer’s weight matrix). This fine-grained allocation is guided by the objective of reducing variance in quality retention across different outputs during the pruning process. We demonstrate that TRIM achieves state-of-the-art pruning performance at extreme sparsity levels (e.g., upwards of 80% of parameters removed).
Because TRIM adapts sparsity at the dimension-wise level, it is inherently compatible with and can augment existing layer-wise allocation strategies. TRIM can be readily integrated into importance-score-based, one-shot pruning algorithms while maintaining computational efficiency. The contributions of our work are as follows:
@@
    We introduce TRIM, the first gradient-free, one-shot pruning algorithm designed to calculate sparsity ratios on a per-dimension basis.
@@
    Using TRIM, we achieve state-of-the-art pruning results at extreme levels of sparsity reducing the perplexity of OPT-13B at 80% from 6461 with Wanda to 324 with TRIM.
@@ 2 Related Work
While these methods achieve strong results, the computational burden of retraining models with billions of parameters has driven the development of efficient one-shot pruning approaches that eliminate the need for retraining. SparseGPT (Frantar and Alistarh, 2023) is the first to apply one-shot pruning to large LLMs. However, to compensate for pruning errors, it requires updating the remaining weights by calculating expensive Hessian inversions. Wanda (Sun et al., 2024) offers a more efficient alternative without updating weights. It addresses the limitations of magnitude pruning (Han et al., 2015) (which fails due to LLMs having outlier activations) by incorporating input statistics of activations. Wanda assigns a score to each weight based on both the absolute weight value and the L2 norm of a set of calibration inputs ($|W|\cdot||X||_{2}$). Further, Wanda innovates the selection of the comparison group - the subset of candidate weights from which the lowest-scoring weights are identified and removed to meet a sparsity target $T$. Wanda adopts a per-output (or row-wise) comparison, pruning the $T\%$ of lowest-scoring weights independently for each output dimension in the weight matrix. This per-output strategy departs from prior approaches, like Magnitude pruning, which utilize the entire layer as the comparison group. However, the same per-output structure that contributes to Wanda’s efficiency can also be a limitation, causing it to struggle at higher sparsity ratios compared to techniques with weight updates like SparseGPT (Frantar and Alistarh, 2023).
Methods such as Outlier Weighed Layerwise (OWL) sparsity extend Wanda by introducing layer-wise sparsity ratios (Yin et al., 2024). OWL assigns unique sparsity ratios to individual layers of a model while maintaining a global sparsity target. By assigning unique sparsity ratios to each layer, OWL avoids pruning layers that contain many outlier positions. An outlier position exists when its Wanda score, $A_{ij}$, is greater than a predefined multiple, $M$, of the mean Wanda score of that layer, $\bar{A}$. Namely, $A_{ij}$ is an outlier position if $A_{ij}>M\cdot\bar{A}$. AlphaPruning (Lu et al., 2024) offers an alternative to OWL by assigning layer-wise sparsity ratios based on the heavy-tailed shape of spectral densities in layer weight matrices. It has its roots in heavy-tailed self-regularization theory (Martin and Mahoney, 2019a, b), making it a more principled approach than OWL.
@@
3 TRIM - Targeted Row-wise Iterative Metric-driven pruning
@@
1:  Input: Weight matrix $W$, Input activations $X$, Target average sparsity $T$, learning rate $\alpha$
2:  Functions: Prune(), Qmetric(), QmetricDimwise()
3:  Output: Optimal dimension-wise sparsity vector $S_{best}$.
4:  Compute unpruned output: $Y\leftarrow WX$
5:  Start with uniform sparsity: $S_{i}\leftarrow T$ for $i=1,\dots,D$
6:  Track best sparsity vector: $S_{best}\leftarrow S$
7:  Track best quality score: $q_{best}\leftarrow-\infty$
8:  for $k=0$ to $K-1$ do
9:     Obtain pruned weights: $W_{pruned}\leftarrow$ Prune(W, S)
10:     Compute pruned output: $\hat{Y}\leftarrow W_{pruned}X$
11:     Calculate overall pruning quality for this iteration: $q_{k}\leftarrow\text{Qmetric}(Y,\hat{Y})$
12:     if $q_{k}>q_{best}$ then
13:        $q_{best}\leftarrow q_{k}$
14:        $S_{best}\leftarrow S$
15:     end if
16:     Calculate quality per-output: $c_{i}\leftarrow\text{QmetricDimwise}(Y_{i,:},\hat{Y}_{i,:})$ for $i=1,\dots,D$
17:     Normalize similarities to [0, 1]: $c^{\prime}_{i}\leftarrow\frac{c_{i}-\min_{j}c_{j}}{\max_{j}c_{j}-\min_{j}c_{j}+\epsilon}$ for $i=1,\dots,D$
18:     Apply learning rate: $\delta_{i}\leftarrow\alpha c^{\prime}_{i}$ for $i=1,\dots,D$
19:     Recenter mean: $S_{i}\leftarrow\delta_{i}-\frac{1}{D}\sum_{j=1}^{D}\delta_{j}+T$ for $i=1,\dots,D$
20:  end for
21:  Return optimal sparsity allocation: Return $S_{best}$
Figure 1 illustrates the core concept of TRIM. Let $W\in\mathbb{R}^{D\times N}$ be the weight matrix of a layer, where $D$ is the number of output dimensions (rows) and $N$ is the number of input dimensions (columns). Each row $W_{i,:}$ corresponds to the weights associated with the $i$-th output dimension. We define a dimension-wise sparsity vector $S=[S_{1},S_{2},\dots,S_{D}]$, where each element $S_{i}\in[0,1]$ specifies the target sparsity ratio for the $i$-th output dimension $W_{i,:}$. This means that for each row $W_{i,:}$, $S_{i}\cdot N$ of its weights will be pruned.
While the individual sparsity ratios $S_{i}$ can vary, they must collectively satisfy the target sparsity $T$ for the layer. This is enforced by ensuring their average equals $T$:
@@
To find a suitable dimension-wise sparsity vector $S$, we employ an iterative adjustment process, detailed in Algorithm 1. The process begins by computing the unpruned output, $Y\leftarrow WX$, (Line 4) with $Y\in\mathbb{R}^{D\times L}$ and initializing $S$ uniformly, such that $S_{i}=T$ for all output dimensions $i=1,\dots,D$ (Line 5). This initial $S$ also serves as the first candidate for the best sparsity vector, $S_{best}$.
The algorithm then iterates $K$ times (Lines 8-20). In each iteration $k$:
@@
    The overall quality $q_{k}\leftarrow\text{Qmetric}(Y,\hat{Y})$ of the iteration is calculated (Line 11). If $q_{k}$ is better than the best quality found so far ($q_{best}$), then $q_{best}$ is updated, and the current $S$ is stored as $S_{best}$ (Lines 12-14).
@@
After $K$ iterations, the algorithm returns $S_{best}$ (Line 21), which represents the dimension-wise sparsity allocation that yields the highest overall quality $q_{best}$ during the iterative search. $S_{best}$ is then used to perform the final pruning of the layer.
To find the most effective sparsity distribution, we first explore positive learning rates, $\alpha$. We start with a small $\alpha$, run the iterative adjustment process (Algorithm 1), and record $q_{best}$. We then incrementally increase $\alpha$ and repeat the process, stopping when a larger $\alpha$ no longer improves $q_{best}$. However, if a positive $\alpha$ does not yield a quality $q_{best}$ superior to the initial baseline (uniform sparsity), we use a negative learning rate. This allows for better generalization across different model activation spaces.
@@
Figure 2: Perplexity progression from 70% to 80% sparsity. TRIM-enhanced OWL shows delayed and slower perplexity growth, effectively extending the usable sparsity frontier.
@@ 4 Experiments
We evaluate TRIM on Qwen2.5 3B/7B/14B/32B/72B (Qwen et al., 2025), LLaMA-2 7B/13B (Touvron et al., 2023), and OPT 7B/13B (Zhang et al., 2022). Our evaluation adheres to established methodologies for pruning LLMs (Sun et al., 2024; Frantar and Alistarh, 2023). We evaluate language modeling proficiency by calculating perplexity on the reserved WikiText validation set (Merity et al., 2016) and assess zero-shot performance across a range of down-stream tasks. We evaluate on BoolQ (Clark et al., 2019), RTE (Wang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC Easy and Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018), using the framework by Gao et al. (Mihaylov et al., 2018).
@@
We evaluate Wanda (Sun et al., 2024) based OWL (Yin et al., 2024) and AlphaPruning (Lu et al., 2024) with and without TRIM. For all runs, we use the same number of calibration samples, each with a sequence length of 2048 tokens, randomly selected from the first shard of the C4 dataset (Raffel et al., 2023). More details and hyperparameters for the layerwise methods can be found in 11.
4.1 Main results
We evaluate pruning at high sparsity ratios of 70% and 80%. For all methods, we focus on pruning the linear layers, excluding the initial embedding layer and the final classification head.
Language Modeling.
Table 1 shows perplexity results for all models used in this study. Integrating TRIM with Wanda, OWL and AlphaPruning consistently results in lower perplexity values than these baselines without TRIM. At 70% and 80% sparsity respectively, TRIM reduces perplexity for LLaMA-2-13B by 10% and 31%, and for Qwen2.5-14B by 9% and 48%. The effectiveness of TRIM is especially pronounced for OPT-13B at 80% sparsity, where TRIM decreases OWL pruning perplexity from 6461.43 to 324.14.
TRIM also improves the robustness and reliability of pruning outcomes. Baseline methods can occasionally lead to counterintuitive scaling results. For example, pruning Qwen2.5-14B at 80% sparsity has a higher perplexity than its 7B counterpart. Baseline methods fail to effectively prune OPT models at 80% sparsity, resulting in very high perplexity. TRIM delivers substantial improvements for OPT stabilizing performance at high sparsity levels. These consistent and robust improvements across all model families, sizes, and sparsity levels underscore the critical role of effective dimension-wise sparsity allocation and highlight TRIM’s efficacy in achieving it.
Zero-Shot Downstream Tasks.
The benefits of TRIM extend beyond perplexity improvements to generalization on downstream tasks. As evidenced by the average zero-shot accuracies presented in Table 2, integrating TRIM with layer-wise pruning strategies yields superior performance compared to baseline methods without TRIM. This performance advantage holds across a wide range of models and sparsity levels. The complete result tables can be found in Appendices 12 and 13.

### v2

@@ Abstract
Large Language Models (LLMs) present significant computational and memory challenges due to their extensive size, making pruning essential for their efficient deployment. Existing one-shot pruning methods often apply uniform sparsity constraints across layers or within each layer, resulting in suboptimal performance at high sparsity ratios. This work introduces TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that applies varying sparsity ratios to individual output dimensions (rows) within each layer. TRIM employs an iterative adjustment process guided by quality metrics to optimize dimension-wise sparsity allocation, ensuring a more equal quality retention across output dimensions to preserve critical information. TRIM can be seamlessly integrated with existing layer-wise pruning strategies. Our evaluations on perplexity and zero-shot tasks across diverse LLM families (Qwen2.5, LLaMA-2, and OPT) and sparsity levels demonstrate that TRIM achieves new state-of-the-art results and enhances stability. For instance, at 80% sparsity, TRIM reduces perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to baseline methods. We conclude that fine-grained, dimension-wise sparsity adaptation is crucial for pushing the limits of extreme LLM compression. Code available at: github.com/flobk/TRIM
TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning
Florentin Beck1, William Rudman2, Carsten Eickhoff1, 1 University of Tübingen, School of Medicine 2 University of Texas at Austin, Department of Linguistics florentin.beck@student.uni-tuebingen.de
@@ 1 Introduction
Large language models (LLMs) have gained prominence due to their generalized problem-solving abilities (Brown et al., 2020; OpenAI et al., 2024; Zhao et al., 2025). A significant driver of this success has been the exponential scaling of model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Increasing the number of parameters has improved LLM performance and enabled emergent behaviors (Wei et al., 2022), such as in-context learning (Brown et al., 2020) and reasoning (Yao et al., 2023; Wei et al., 2023). However, this growth in model size has lead to substantial memory and computational demands, which pose considerable deployment challenges (Zhou et al., 2024).
@@
As a result, one-shot pruning methods that reduce LLM parameters without retraining to compute weight masks have gained prominence (Han et al., 2015; Sun et al., 2024; Zhang et al., 2024a; Das et al., 2024). A challenge with current approaches is that they often impose predefined structural constraints, such as uniform sparsity across all network layers or outputs. This imposed structure is arbitrary and can be detrimental to pruning performance, particularly as LLMs exhibit unique weight and activation characteristics, such as prominent outlier features and highly skewed activation distributions (Kovaleva et al., 2021; Dettmers et al., 2022; Rudman et al., 2024).
Layer-wise sparsity adaptation methods aim to mitigate this by assigning distinct sparsity ratios to each layer of the model, determined through algorithmic search (Li et al., 2024; Xu et al., 2025), heuristic-based approaches (Yin et al., 2024; Sun et al., 2025), or theoretically grounded strategies (Lu et al., 2024). While this offers more flexibility than uniform sparsity, these methods still apply a consistent sparsity target within each layer and can struggle under extreme sparsity levels, such as removing more than 70% of model parameters.
In this work, we propose TRIM: (Targeted Row-wise Iterative Metric-driven pruning). TRIM builds on the principles of one-shot pruning algorithms like Wanda (Sun et al., 2024), which compare and remove weights at each layer iteratively. Current methods typically apply a uniform sparsity constraint across all output dimensions (per-output), however, TRIM introduces a more granular approach. TRIM calculates unique sparsity ratios for individual output dimensions (corresponding to a row in the layer’s weight matrix). This fine-grained allocation is guided by the objective of reducing variance in quality retention across different outputs during the pruning process. We demonstrate that TRIM achieves state-of-the-art pruning performance at extreme sparsity levels (e.g., upwards of 80% of parameters removed).
Because TRIM adapts sparsity at the dimension-wise level, it is inherently compatible with and can augment existing layer-wise allocation strategies. In this paper, we use TRIM to improve two layer-wise pruning methods: OWL and AlphaPruning.
TRIM readily integrates into importance-score-based pruning algorithms, providing a principled extension of such methods beyond any single metric (e.g., not limited to Wanda). In doing so, TRIM preserves computational efficiency and broadens the applicability of per-output pruning approaches. The contributions of our work are as follows:
@@
    We introduce TRIM, the first algorithm to enhance pruning granularity by calculating sparsity ratios on a per-dimension basis.
@@
    Using TRIM, we preserve performance at lower sparsity ratios and achieve state-of-the-art pruning results at extreme levels of sparsity reducing the perplexity of OPT-13B at 80% from 6461 with Wanda-based OWL to 324 with TRIM.
@@ 2 Related Work
While these methods achieve strong results, the computational burden of retraining models with billions of parameters has driven the development of efficient one-shot pruning approaches that eliminate the need for retraining. Zhang et al. (2024b); van der Ouderaa et al. (2024)
[…]
Methods such as Outlier Weighed Layerwise (OWL) sparsity extend Wanda by introducing layer-wise sparsity ratios (Yin et al., 2024). OWL assigns unique sparsity ratios to individual layers of a model while maintaining a global sparsity target. By assigning unique sparsity ratios to each layer, OWL avoids pruning layers that contain many outlier positions. An outlier position exists when its Wanda score, $A_{ij}$, is greater than a predefined multiple, $M$, of the mean Wanda score of that layer, $\bar{A}$. Namely, if $A_{ij}>M\cdot\bar{A}$. AlphaPruning (Lu et al., 2024) offers an alternative to OWL by assigning layer-wise sparsity ratios based on the heavy-tailed shape of spectral densities in layer weight matrices. It has its roots in heavy-tailed self-regularization theory (Martin and Mahoney, 2019a, b), making it a more principled approach than OWL.
@@
[…]
3 TRIM
Figure 1 illustrates the core concept of TRIM. Let $W\in\mathbb{R}^{D\times N}$ be the weight matrix of a layer, where $D$ is the number of output dimensions (rows) and $N$ is the number of input dimensions (columns). Each row $W_{i,:}$ corresponds to the weights associated with the $i$-th output dimension. We define a dimension-wise sparsity vector $S=[S_{1},S_{2},\dots,S_{D}]$, where each element $S_{i}\in[0,1]$ specifies the target sparsity ratio for the $i$-th output dimension $W_{i,:}$. This means that for each row $W_{i,:}$, $S_{i}\cdot N$ of its weights will be pruned. While the individual sparsity ratios $S_{i}$ can vary, they must collectively satisfy the target sparsity $T$ for the layer. This is enforced by ensuring their average equals $T$:
@@ 3 TRIM
To find a suitable dimension-wise sparsity vector $S$, we employ an iterative adjustment process detailed in Algorithm 1. The process begins by computing the unpruned output $Y\leftarrow WX$ (Line 4), with $Y\in\mathbb{R}^{D\times L}$ and initializing $S$ uniformly, so that $S_{i}=T$ for all output dimensions $i=1,\dots,D$ (Line 5). This initial $S$ also serves as the first candidate for the best sparsity vector, $S_{best}$. The algorithm then iterates $K$ times (Lines 8-20). In each iteration $k$:
@@
[…]
@@
After $K$ iterations, the algorithm returns $S_{best}$ (Line 21), which represents the dimension-wise sparsity allocation that yields the highest overall quality $q_{best}$ during the iterative search. $S_{best}$ is then used to perform the final pruning of the layer. In Appendix G, we explore the impact of #$k$ on the pruning quality of TRIM.
To determine the optimal sparsity distribution, we evaluate positive learning rates $\alpha$, beginning with small values, applying Algorithm 1, and recording $q_{best}$. We increase $\alpha$ and repeat the process, stopping when a larger $\alpha$ no longer improves $q_{best}$. However, if a positive $\alpha$ does not yield a quality $q_{best}$ superior to the uniform sparsity baseline, we use a negative learning rate. This allows for better generalization across different model activation spaces. In Section 5, we briefly explore when a negative learning rate occurs in models.
@@
Figure 2: Perplexity progression from 70% to 80% sparsity. TRIM extends the usable sparsity range.
@@ 4 Experiments
We evaluate TRIM on Qwen2.5 3B/7B/14B/32B/72B (Qwen et al., 2025), LLaMA-2 7B/13B (Touvron et al., 2023), and OPT 7B/13B (Zhang et al., 2022). Our evaluation adheres to established methodologies for pruning LLMs (Sun et al., 2024; Frantar and Alistarh, 2023). We evaluate language modeling proficiency by calculating perplexity on the reserved WikiText validation set (Merity et al., 2016) and assess zero-shot performance across a range of down-stream tasks. We evaluate on BoolQ (Clark et al., 2019), RTE (Wang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC Easy and Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018), using the framework by Gao et al. (Mihaylov et al., 2018). Additionally, we evaluate the generalizability of the WikiText results on C4 and Pile in Appendix F.
@@
We evaluate Wanda (Sun et al., 2024) based OWL (Yin et al., 2024) and AlphaPruning (Lu et al., 2024) with and without TRIM. In Appendix C, we evaluate Wanda with uniform layerwise sparsities. For all runs, we use the same number of calibration samples, each with a sequence length of 2048 tokens, randomly selected from the first shard of the C4 dataset (Raffel et al., 2023). Hyperparameters can be found in Appendix H.
@@
4.1 Main results
We evaluate OWL and AlphaPruning with and without TRIM at high sparsity ratios (70% and 80%), and additionally evaluate OWL with and without TRIM at 50% and 60%. For all methods, we only prune linear layers, excluding the initial embedding layer and the final classification head.
Language Modeling.
Table 1 shows perplexity results for all models used in this study. Integrating TRIM with OWL and AlphaPruning consistently results in lower perplexity values than these baselines without TRIM. At 70% and 80% sparsity respectively, TRIM reduces perplexity for LLaMA-2-13B by 10% and 31%, and for Qwen2.5-14B by 9% and 48%. The effectiveness of TRIM is especially pronounced for OPT-13B at 80% sparsity, where TRIM decreases OWL pruning perplexity from 6461.43 to 324.14.
TRIM improves the robustness and reliability of pruning results, correcting counterintuitive behaviors seen with baseline methods. At 80% sparsity, Qwen2.5-14B shows higher perplexity than its smaller 7B counterpart, and OPT models suffer from unusually poor performance. TRIM resolves both issues, restoring expected scaling and stable performance. These consistent and robust improvements across all model families, sizes, and sparsity levels underscore the critical role of effective dimension-wise sparsity allocation and highlight TRIM’s efficacy in achieving it.
Table 3: Perplexity results on WikiText validation set for various models pruned at 50% and 60% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 50% OWL 12.39 12.10 7.24 6.09 12.04 8.80 8.16 6.31 5.29 -0.8% +TRIM 12.39 12.08 7.15 5.99 11.94 8.79 8.12 6.32 5.27 60% OWL 16.12 16.39 9.65 7.82 24.91 9.65 11.67 7.64 6.70 -2.3% +TRIM 16.04 16.38 9.36 7.49 23.53 9.35 11.36 7.58 6.63
Zero-Shot Downstream Tasks.
The benefits of TRIM extend beyond perplexity improvements to generalization on downstream tasks. As evidenced by the average zero-shot accuracies presented in Table 2, integrating TRIM with layer-wise pruning strategies yields superior performance compared to baseline methods without TRIM. This performance advantage holds across a wide range of models and sparsity levels. The complete result tables can be found in Appendices 16 and 17.
4.2 Lower Sparsity Levels.
At moderate sparsity (50-60%), TRIM matches or improves WikiText validation perplexity relative to the OWL baseline across all tested models. These results (Table 3) indicate that TRIM’s benefits persist below the high-sparsity regime, enhancing or preserving quality also at lower pruning ratios.